# Detection of pairwise kSZ effect with DESI galaxy groups and Planck in Fourier space

### 07 Jan 2024 v1

@@ 6 Conclusion and discussion
In this work, we report a $\sim 5.2\sigma$ detection of the kinetic Sunyaev–Zel’dovich (kSZ) effect in Fourier space, by combining the DESI galaxy groups and the Planck data. By implying the pairwise estimator, we measure the density-weighted pairwise kSZ power spectrum of the galaxy group sample, together with the Planck HFI 217 GHz data. The galaxy group catalog is constructed from the DR9 data of the DESI Legacy Imaging Surveys, and is selected by varying the lower limit of the group mass, which is estimated from the group luminosity (Yang et al., 2021). The baseline sample we construct has a lower mass limit of $10^{13}{\rm M}_{\odot}/h$ and a median mass $\tilde{M}=1.75\times 10^{13}{\rm M}_{\odot}/h$. It gives the highest S/N$=5.2$ when the adopted AP filter radius $\theta_{\rm AP}=4.2^{\prime}$. By fitting $\bar{\tau}$s from samples with different median $M$, we obtain a linear $\log\bar{\tau}=\gamma\log(\tilde{M}-14)+\log\beta$, in which $\gamma=0.55^{+0.1}_{-0.1}$.
@@
The optical depth $\bar{\tau}$ is constrained by implementing a joint analysis of kSZ and RSD effects on the sample group catalog. The RSD analysis determines the group bias parameters $b_{1}$, $b_{2}$ and the parameter $\sigma_{z}$ quantifying the photo-z damping effect. Then the best fitted three parameters are substituted to the kSZ analysis and constrain the optial depth $\bar{\tau}$. The stability of the $\bar{\tau}$ measurement against variations on the photo-z damping functional form and the prior knowledge of $b_{1}$ from the $b_{1}-M$ model is verified in figure 11. Although the degeneracy exists between $\bar{\tau}$ and $b_{1}$, and between $\bar{\tau}$ and the photo-z damping effect, it does not induce significant bias of the measured $\bar{\tau}$ in this work. Alternatively speaking, in a joint kSZ and RSD analysis, the measured $\hat{P}_{\ell}$ determines the shape and amplitude of $P_{\rm pv}^{\ell}$ and in turn the corresponding $\bar{\tau}$. The details about the $b_{1}$, $b_{2}$ and $\sigma_{z}$ fitting results, including how they depend on the adopted photo-z damping functional form, is not so important for the purpose of a $\bar{\tau}$ measurement. Indeed, the joint kSZ and RSD analysis in this work is not rigorous, in that we do not apply a global fitting of measured $\hat{P}_{\ell}$ and $\hat{P}^{\ell}_{\rm kSZ}$ simultaneously. In the future, this joint fitting of kSZ and RSD effects can serve as a standard practice of the large scale structure analysis, by which we can constrain the cosmic structure growth and the baryonic distribution in and around halos at the same time (Sugiyama et al., 2017; Okumura & Taruya, 2022).
@@
The kSZ effect is a good indicator of the baryonic distribution in the universe. By varying the AP filter radius, we also measure the $\bar{\tau}$ profiles of the group samples, which is tightly coupled with the baryon density profiles within halos. In a companion paper, we will study these $\bar{\tau}$ profiles with more adequate gas density profile models, e.g., the baryonification models (Schneider & Teyssier, 2015; Giri & Schneider, 2021; Aricò et al., 2021), and constrain the associated baryon density profiles (Schneider et al., 2022). In turn, we can constrain the baryonic damping of the dark matter density power spectrum, which is a key systematic error of the weak lensing cosmology (Jing et al., 2006; Schneider et al., 2022; Chen et al., 2023a).
@@
With the next generation of CMB experiments and galaxy redshift surveys, the detection S/N of the kSZ effect can reach $\mathcal{O}(100)$ and beyond (Sugiyama et al., 2018a; Smith et al., 2018). At that time, the kSZ effect can be robustly used to help measure the cosmic structure growth history and constrain cosmological models (Sugiyama et al., 2017; Zheng, 2020; Okumura & Taruya, 2022). In this sense, the Fourier space kSZ analysis will has its unique advantage on its easier redshift space modelling and the covariance matrix evaluation. This work can be regarded a necessary preparation of this exciting era.
@@
Acknowledgement: We thank Pengjie Zhang, Teppei Okumura and Kwan Chuen Chan for useful discussions. We thank the referee Naonori Sugiyama for bringing up valuable comments and heavily improve the paper quality. YZ acknowledges the supports from the National Natural Science Foundation of China (NFSC) through grant 12203107, the Guangdong Basic and Applied Basic Research Foundation with No.2019A1515111098, and the science research grants from the China Manned Space Project with NO.CMS-CSST-2021-A02. XHY acknowledges the supports from National Key R&D Program of China (2023YFA1607800, 2023YFA1607804).
@@ Appendix A RSD analysis of the DESI photo-z groups
In this appendix we measure the redshift space density power spectrum multipoles $\hat{P}_{\ell}$ of the group sample. By fitting eqs. (11) and (16) to the measured $\hat{P}_{\ell}$, we determine the posterior distributions of the bias parameters $b_{1}$, $b_{2}$ and the photo-z error dispersion $\sigma_{z}$. The best fitted values of these parameters are then substituted into eq. (14) to calculate the non-linear kSZ power spectrum and fit the $\bar{\tau}$ parameter in the main content of this paper.
@@ A.1 $\hat{P}_{\ell}$ measurements
where the normalization factor $A$ is calculated in the same way as that of $\hat{P}^{\ell}_{\rm kSZ}$ in eq. (27), and the shot noise term is
@@
In consistency with $\hat{P}^{\ell}_{\rm kSZ}$, we do not apply the FKP weight in measuring $\hat{P}_{\ell}$. We do not apply the imaging systematic weights here as well, since its impact on the $\hat{P}_{\ell}$ measurement is negligible as demonstrated in section 4.8.
The measured baseline sample $\hat{P}_{\ell}$s are shown on the left panel of figure 9. Compared with $\hat{P}_{\ell}$ of a spec-z sample, e.g. figure 6 of Beutler et al. (2017), we see a huge photo-z damping effect. The photo-z error of $\sigma_{\rm pho-z}\sim\mathcal{O}(0.01)$ corresponds to a $\sigma_{z}\sim\mathcal{O}(3000){\rm km/s}$. It heavily damps the measured monopole $\hat{P}_{0}$ even at linear scales. This damping naturally generated a degeneracy between $b_{1}$ and $\sigma_{z}$, which in turn affects the $\bar{\tau}$ measurement. We will test the potential bias of the $\bar{\tau}$ measurement due to an improper treatment of the photo-z uncertainty in this appendix. Moreover, for a spec-z sample, $\hat{P}_{2}$ is positive and $\hat{P}_{4}$ is nearly zero, while in figure 9 $\hat{P}_{2}$ becomes negative and $\hat{P}_{4}$ is away from zero in a significant level. The photo-z uncertainty leaves significant imprints on $\hat{P}_{2}$ and $\hat{P}_{4}$, from which we can potentially infer its properties.
@@ A.2 Fitting results
Figure 10: Left: Window function multipoles evaluated from the random catalog corresponding to the NGC sub-baseline sample . Right: Theoretical masked and unmasked power spectrum multipoles and their fractional differences.
Figure 11: Left: MCMC fitting results of multipoles of power spectrum of our baseline sample with three free parameters. Different colors indicate results of different photo-z damping functions. Upper Right: The $b_{1}-\tilde{M}$ relationship, where $b_{1}$s are obtained by MCMC fittings (data points with error bars) or given by theoretical models (solid lines). Lower Right: Measured $\bar{\tau}-\tilde{M}$ relationships when different $D_{\rm pho-z}$s and $b_{1}-M$ models are adopted. The black data points with error bars are results when we fit $b_{1}$, $b_{2}$ and $\sigma_{z}$ simultaneously. The colored points are results when we first calculate $b_{1}$ from the corresponding $b_{1}-M$ model and then fit $b_{2}$ and $\sigma_{z}$ from $\hat{P}_{\ell}$ measurements. The adopted $b_{1}-M$ models embedded in the colossus python toolkit (Diemer, 2018) are: jing98 (Jing, 1998), sheth01 (Sheth et al., 2001), tinker10 (Tinker et al., 2010) and bhattacharya11 (Bhattacharya et al., 2011).
During the $P_{\ell}$ fitting, we include the window function effect in the RSD model, as described in Wilson et al. (2017). As an example, the window function multipoles of the NGC sub-baseline sample are shown on the left panel of figure 10, and differences between the masked and unmasked $P_{\ell}$ models are plotted on the right. As shown, it is necessary to include the window function effect in the model since there are up to $7\%$ differences between models with and without masks at the fitting scales $0.05h/{\rm Mpc}\lesssim k\lesssim 0.15h$/Mpc.
@@
The PDF of the photo-z uncertainty is a complicated, possibly asymmetric function in its nature. Modelling its impact on the redshift space galaxy power spectrum is a difficult topic. In particular, improper modelling of the photo-z damping can lead to systematic errors of the bias and $\sigma_{z}$ parameters. Here we test this possibility by trying three terms of photo-z damping functions, which sequentially correspond to a Gaussian photo-z error PDF, a Lorentzian photo-z error PDF and a Lorentzian pairwise photo-z error PDF (Davis & Peebles, 1983; Ballinger et al., 1996):
@@
All PDFs are assumed to be symmetric. Different lines on the left panel of figure 9 are the best fitted RSD models incorporating the above damping functions. Visually $P_{\rm s}^{\rm pho-z}$ with the Lorentzian $D_{\rm pho-z}$ best fits the data. Therefore in this work we will mainly adopt the Lorentzian damping term in modelling $P_{\rm s}^{\rm pho-z}$ and $P_{\rm kSZ}^{\rm pho-z}$.
@@
We can compare the fitted $b_{1}$ and $\sigma_{z}$ with some prior knowledge we have. As discussed in section 2.2, we expect that the photo-z uncertainty $\sigma_{\rm pho-z}$ of our samples to be $0.005\sim 0.025$. It corresponds to $\sigma_{z}\equiv c\sigma_{\rm pho-z}=0.15\times 10^{4}\sim 0.75\times 10^{4}$km/s. This expectation is closer to the fitted $\sigma_{z}$ of the Gaussian photo-z damping function, and lower than those of the Lorentzian2 and Lorentzian functions. On the upper right panel of figure 11, we compare the fitted $b_{1}$s with predictions from several $b_{1}-M$ models in the literature. Within three $D_{\rm pho-z}$ models, the $b_{1}-\sigma_{z}$ degeneracy indicates a higher $b_{1}$ from a higher fitted $\sigma_{z}$. Moreover, the fitted $b_{1}$s of low $\tilde{M}$ samples are comparable to model predictions, while for high $\tilde{M}$ samples, our fitting results are higher than model predictions.
These discrepancies, although not necessarily mean that any of the fitted $b_{1}$ and $\sigma_{z}$ are definitely wrong, probably indicate an inadequate model of the photo-z damping effect. Luckily, the main target of this work is to determine an accurate $\bar{\tau}$, with the accuracy of the fitted $b_{1}$ and $\sigma_{z}$ being less significant. On the lower right panel of figure 11, we compare fitted $\bar{\tau}$s of different samples when various photo-z damping functions and $b_{1}-M$ models are adopted. It is shown that fitted $\bar{\tau}$s are quite insensitive to these model uncertainties, so we are confident that main results of this work are immune from our poor understanding of the photo-z uncertainty.
@@ Appendix B Convergence test of the $Q_{\ell}$ estimation
Figure 12: Left: Window function multipoles evaluated from the random catalog corresponding to the NGC sub-baseline sample. We use the random samples whose sizes are equal to 2, 1 and 0.5 times of the group sample size $N_{\rm group}$. The lower panel shows the fractional difference between those of “$2\times N_{\rm group}$” and “$1\times N_{\rm group}$” (blue lines) or “$0.5\times N_{\rm group}$” (green lines). Right: Theoretical masked dipoles and their fractional differences.
In order to speed up the window function calculation, for samples with $M_{\rm th}\geq 10^{13}{\rm M}_{\odot}/h$, we randomly dilute the random catalog to the size of the $M_{\rm th}=10^{13}{\rm M}_{\odot}/h$ group sample, and for samples with $M_{\rm th}<10^{13}{\rm M}_{\odot}/h$, we dilute the random catalog to the size of the corresponding group sample size. On the left panel of figure 12, we compare the estimated $Q_{0}$ and $Q_{2}$ of the NGC sub-baseline sample by using 1/10, 1/20 and 1/40 of all random points, corresponding to a random sample size of 2, 1 and 0.5 times the group sample size. We observe good convergence of $Q_{\ell}$, in particular that for $Q_{0}$, no systematic offsets are presented. On the right panel of figure 12, we compare the masked theoretical dipole evaluated by eq. (43) using different sizes of random samples, and show that the dilution induces little systematic errors in the $\tilde{P}_{\rm kSZ}^{\ell=1}$ calculation.
Within the $k$ range of our interest, similar tests on other group samples give at most $1\%$ systematic offset of the $\tilde{P}_{\rm kSZ}^{\ell=1}$ calculation for the $M_{\rm th}>10^{14}{\rm M}_{\odot}/h$ group sample, which is well within the statistical error of $\bar{\tau}$ estimation and will not affect our conclusions.

### 01 Mar 2024 v2

@@ 6 Conclusion and discussion
In this work, we report a $\sim 5.2\sigma$ detection of the kSZ effect in Fourier space, by combining the DESI galaxy groups and the Planck data. By implying the pairwise estimator, we measure the density-weighted pairwise kSZ power spectrum of the galaxy group sample, together with the Planck HFI 217 GHz data. The galaxy group catalog is constructed from the DR9 data of the DESI Legacy Imaging Surveys, and is selected by varying the lower limit of the group mass, which is estimated from the group luminosity (Yang et al., 2021). The baseline sample we construct has a lower mass limit of $10^{13}M_{\odot}/h$ and a median mass $\tilde{M}=1.75\times 10^{13}M_{\odot}/h$. It gives the highest S/N$=5.2$ when the adopted AP filter radius $\theta_{\rm AP}=4.2^{\prime}$. By fitting $\bar{\tau}$s from samples with different median $M$, we obtain a linear $\log\bar{\tau}=\gamma\log(\tilde{M}-14)+\log\beta$, in which $\gamma=0.55^{+0.1}_{-0.1}$.
@@
The optical depth $\bar{\tau}$ is constrained by implementing a joint analysis of kSZ and RSD effects on the sample group catalog. The RSD analysis determines the group bias parameters $b_{1}$, $b_{2}$ and the parameter $\sigma_{z}$ quantifying the photo-z damping effect. Then, the best-fitted three parameters are substituted to the kSZ analysis and constrain the optial depth $\bar{\tau}$. The stability of the $\bar{\tau}$ measurement against variations on the photo-z damping functional form and the prior knowledge of $b_{1}$ from the $b_{1}-M$ model is verified in figure 11. Although the degeneracy exists between $\bar{\tau}$ and $b_{1}$, and between $\bar{\tau}$ and the photo-z damping effect, it does not induce significant bias of the measured $\bar{\tau}$ in this work. Alternatively speaking, in a joint kSZ and RSD analysis, the measured $\hat{P}_{\ell}$ determines the shape and amplitude of $P_{\rm pv}^{\ell}$ and in turn the corresponding $\bar{\tau}$. The details about the $b_{1}$, $b_{2}$ and $\sigma_{z}$ fitting results, including how they depend on the adopted photo-z damping functional form, are not so important for the purpose of a $\bar{\tau}$ measurement. Indeed, the joint kSZ and RSD analysis in this work is not rigorous, in that we do not apply a global fitting of measured $\hat{P}_{\ell}$ and $\hat{P}^{\ell}_{\rm kSZ}$ simultaneously. In the future, this joint fitting of kSZ and RSD effects can serve as a standard practice of the large-scale structure analysis, by which we can constrain the cosmic structure growth and the baryonic distribution in and around halos at the same time (Sugiyama et al., 2017; Okumura & Taruya, 2022).
@@
The kSZ effect is a good indicator of the baryonic distribution in the universe. By varying the AP filter radius, we also measure the $\bar{\tau}$ profiles of the group samples, which is tightly coupled with the baryon density profiles within halos. In a companion paper, we will study these $\bar{\tau}$ profiles with more adequate gas density profile models, e.g., the baryonification models (Schneider & Teyssier, 2015; Giri & Schneider, 2021; Aricò et al., 2021), and constrain the associated baryon density profiles (Schneider et al., 2022). In turn, we can constrain the baryonic damping of the dark matter density power spectrum, which is a key systematic error of the weak-lensing cosmology (Jing et al., 2006; Schneider et al., 2022; Chen et al., 2023a; Zheng & Zhang, 2024).
@@
With the next generation of CMB experiments and galaxy redshift surveys, the detection S/N of the kSZ effect can reach $\mathcal{O}(100)$ and beyond (Sugiyama et al., 2018a; Smith et al., 2018). At that time, the kSZ effect can be robustly used to help measure the cosmic structure growth history and constrain cosmological models (Sugiyama et al., 2017; Zheng, 2020; Okumura & Taruya, 2022). In this sense, the Fourier space kSZ analysis will has its unique advantage on its easier redshift space modeling and the covariance matrix evaluation. This work can be regarded a necessary preparation of this exciting era.
@@
Acknowledgement: We thank Pengjie Zhang, Teppei Okumura, and Kwan Chuen Chan for useful discussions. We thank the referee Naonori Sugiyama for bringing up valuable comments and heavily improving the paper quality. Y.Z. acknowledges the supports from the National Natural Science Foundation of China (NFSC) through grant 12203107, the Guangdong Basic and Applied Basic Research Foundation with No.2019A1515111098, and the science research grants from the China Manned Space Project with NO.CMS-CSST-2021-A02. X.Y. acknowledges the supports from National Key R&D Program of China (2023YFA1607800, 2023YFA1607804).
@@ Appendix A RSD analysis of the DESI photo-z groups
In this appendix, we measure the redshift space density power spectrum multipoles $\hat{P}_{\ell}$ of the group sample. By fitting Equations (11) and (16) to the measured $\hat{P}_{\ell}$, we determine the posterior distributions of the bias parameters $b_{1}$, $b_{2}$ and the photo-z error dispersion $\sigma_{z}$. The best-fitted values of these parameters are then substituted into Equation (14) to calculate the nonlinear kSZ power spectrum and fit the $\bar{\tau}$ parameter in the main content of this paper.
@@ A.1 $\hat{P}_{\ell}$ measurements
where the normalization factor $A$ is calculated in the same way as that of $\hat{P}^{\ell}_{\rm kSZ}$ in Equation (27), and the shot noise term is
@@
In consistency with $\hat{P}^{\ell}_{\rm kSZ}$, we do not apply the FKP weight in measuring $\hat{P}_{\ell}$. We do not apply the imaging-systematic weights here as well, since its impact on the $\hat{P}_{\ell}$ measurement is negligible as demonstrated in Section 4.8.
The measured baseline sample $\hat{P}_{\ell}$s are shown on the left panel of figure 9. Compared with $\hat{P}_{\ell}$ of a spec-z sample, e.g., figure 6 of Beutler et al. (2017), we see a huge photo-z damping effect. The photo-z error of $\sigma_{\rm pho-z}\sim\mathcal{O}(0.01)$ corresponds to a $\sigma_{z}\sim\mathcal{O}(3000){\rm km/s}$. It heavily damps the measured monopole $\hat{P}_{0}$ even at linear scales. This damping naturally generated a degeneracy between $b_{1}$ and $\sigma_{z}$, which in turn affects the $\bar{\tau}$ measurement. We will test the potential bias of the $\bar{\tau}$ measurement due to an improper treatment of the photo-z uncertainty in this appendix. Moreover, for a spec-z sample, $\hat{P}_{2}$ is positive, and $\hat{P}_{4}$ is nearly zero, while in figure 9 $\hat{P}_{2}$ becomes negative, and $\hat{P}_{4}$ is away from zero in a significant level. The photo-z uncertainty leaves significant imprints on $\hat{P}_{2}$ and $\hat{P}_{4}$, from which we can potentially infer its properties.
@@ A.2 Fitting results
Figure 10: Left: window function multipoles evaluated from the random catalog corresponding to the NGC subbaseline sample . Right: theoretical masked and unmasked power spectrum multipoles and their fractional differences.
Figure 11: Left: Markov Chain Monte Carlo (MCMC) fitting results of multipoles of power spectrum of our baseline sample with three free parameters. Different colors indicate results of different photo-z damping functions. Upper Right: the $b_{1}-\tilde{M}$ relationship, where $b_{1}$s are obtained by MCMC fittings (data points with error bars) or given by theoretical models (solid lines). Lower Right: Measured $\bar{\tau}-\tilde{M}$ relationships when different $D_{\rm pho-z}$s and $b_{1}-M$ models are adopted. The black data points with error bars are results when we fit $b_{1}$, $b_{2}$, and $\sigma_{z}$ simultaneously. The colored points are results when we first calculate $b_{1}$ from the corresponding $b_{1}-M$ model and then fit $b_{2}$ and $\sigma_{z}$ from $\hat{P}_{\ell}$ measurements. The adopted $b_{1}-M$ models embedded in the colossus Python toolkit (Diemer, 2018) are: jing98 (Jing, 1998), sheth01 (Sheth et al., 2001), tinker10 (Tinker et al., 2010) and bhattacharya11 (Bhattacharya et al., 2011).
During the $P_{\ell}$ fitting, we include the window function effect in the RSD model, as described in Wilson et al. (2017). As an example, the window function multipoles of the NGC subbaseline sample are shown on the left panel of figure 10, and differences between the masked and unmasked $P_{\ell}$ models are plotted on the right. As shown, it is necessary to include the window function effect in the model since there are up to $7\%$ differences between models with and without masks at the fitting scales $0.05h/{\rm Mpc}\lesssim k\lesssim 0.15h$/Mpc.
@@
The PDF of the photo-z uncertainty is a complicated, possibly asymmetric function in its nature. Modeling its impact on the redshift space galaxy power spectrum is a difficult topic. In particular, improper modeling of the photo-z damping can lead to systematic errors of the bias and $\sigma_{z}$ parameters. Here, we test this possibility by trying three terms of photo-z damping functions, which sequentially correspond to a Gaussian photo-z error PDF, a Lorentzian photo-z error PDF and a Lorentzian pairwise photo-z error PDF (Davis & Peebles, 1983; Ballinger et al., 1996):
@@
All PDFs are assumed to be symmetric. Different lines on the left panel of figure 9 are the best-fitted RSD models incorporating the above damping functions. Visually, $P_{\rm s}^{\rm pho-z}$ with the Lorentzian $D_{\rm pho-z}$ best fits the data. Therefore, in this work we will mainly adopt the Lorentzian damping term in modeling $P_{\rm s}^{\rm pho-z}$ and $P_{\rm kSZ}^{\rm pho-z}$.
@@
We can compare the fitted $b_{1}$ and $\sigma_{z}$ with some prior knowledge we have. As discussed in Section 2.2, we expect that the photo-z uncertainty $\sigma_{\rm pho-z}$ of our samples to be $0.005\sim 0.025$. It corresponds to $\sigma_{z}\equiv c\sigma_{\rm pho-z}=0.15\times 10^{4}\sim 0.75\times 10^{4}$km/s. This expectation is closer to the fitted $\sigma_{z}$ of the Gaussian photo-z damping function, and lower than those of the Lorentzian2 and Lorentzian functions. On the upper right panel of figure 11, we compare the fitted $b_{1}$s with predictions from several $b_{1}-M$ models in the literature. Within three $D_{\rm pho-z}$ models, the $b_{1}-\sigma_{z}$ degeneracy indicates a higher $b_{1}$ from a higher fitted $\sigma_{z}$. Moreover, the fitted $b_{1}$s of low $\tilde{M}$ samples are comparable to model predictions, while, for high $\tilde{M}$ samples, our fitting results are higher than model predictions.
These discrepancies, although not necessarily meaning that any of the fitted $b_{1}$ and $\sigma_{z}$ are definitely wrong, probably indicate an inadequate model of the photo-z damping effect. Luckily, the main target of this work is to determine an accurate $\bar{\tau}$, with the accuracy of the fitted $b_{1}$ and $\sigma_{z}$ being less significant. On the lower right panel of figure 11, we compare fitted $\bar{\tau}$s of different samples when various photo-z damping functions and $b_{1}-M$ models are adopted. It is shown that fitted $\bar{\tau}$s are quite insensitive to these model uncertainties, so we are confident that the main results of this work are immune from our poor understanding of the photo-z uncertainty.
@@ Appendix B Convergence test of the $Q_{\ell}$ estimation
Figure 12: Left: Window function multipoles evaluated from the random catalog corresponding to the NGC subbaseline sample. We use the random samples whose sizes are equal to 2, 1, and 0.5 times of the group sample size $N_{\rm group}$. The lower panel shows the fractional difference between those of “$2\times N_{\rm group}$” and “$1\times N_{\rm group}$” (blue lines) or “$0.5\times N_{\rm group}$” (green lines). Right: theoretical masked dipoles and their fractional differences.
In order to speed up the window function calculation, for samples with $M_{\rm th}\geq 10^{13}M_{\odot}/h$, we randomly dilute the random catalog to the size of the $M_{\rm th}=10^{13}M_{\odot}/h$ group sample, and for samples with $M_{\rm th}<10^{13}M_{\odot}/h$, we dilute the random catalog to the size of the corresponding group sample size. On the left panel of figure 12, we compare the estimated $Q_{0}$ and $Q_{2}$ of the NGC subbaseline sample by using 1/10, 1/20, and 1/40 of all random points, corresponding to a random sample size of 2, 1, and 0.5 times the group sample size. We observe good convergence of $Q_{\ell}$, in particular that for $Q_{0}$, no systematic offsets are presented. On the right panel of figure 12, we compare the masked theoretical dipole evaluated by Equation (43) using different sizes of random samples, and show that the dilution induces little systematic errors in the $\tilde{P}_{\rm kSZ}^{\ell=1}$ calculation.
Within the $k$ range of our interest, similar tests on other group samples give at most $1\%$ systematic offset of the $\tilde{P}_{\rm kSZ}^{\ell=1}$ calculation for the $M_{\rm th}>10^{14}M_{\odot}/h$ group sample, which is well within the statistical error of $\bar{\tau}$ estimation and will not affect our conclusions.